# Christopher Tosh Philip Greengard Ben Goodrich Andrew Gelman Aki Vehtari Daniel Hsu

### v4

@@
The piranha problem: Large effects swimming in a small pond [^1] We thank Lauren Kennedy, the editor, and four anonymous reviewers for helpful discussions. CT and DH acknowledge support from NSF grant CCF-1740833, a JP Morgan Faculty Award, and a Sloan Research Fellowship. Much of this work was done while CT was at Columbia University. AG acknowledges support from ONR grant N000142212648. Author affiliations: Memorial Sloan Kettering Cancer Center (CT), Department of Statistics, Columbia University (PG and AG), Department of Political Science, Columbia University (BG and AG), Department of Computer Science, Aalto University (AV), Department of Computer Science, Columbia University (DH).
@@ 1 Background
The implication of the published and well-publicized claims regarding ovulation and voting, shark attacks and voting, college football and voting, etc., is not merely that some voters are superficial and fickle. No, this literature claims that seemingly trivial or irrelevant factors have large and consistent effects, and this runs into the problem of interactions. For example, the effect on your vote of the local college football team losing could depend crucially on whether there’s been a shark attack lately, or on what’s up with your hormones on election day. Or the effect could be positive in an election with a female candidate and negative in an election with a male candidate. Or the effect could interact with your parents’ socioeconomic status, or whether your child is a boy or a girl, or the latest campaign ad, or any of the many other factors that have been studied in the evolutionary psychology and political psychology literatures. If such effects are large, it is necessary to consider their interactions, because the average effect of a factor in any particular study will depend on the levels of all the other factors in that environment. Similarly, Mellon (2020) has argued against naive assumptions of causal identification in economics, where there is a large literature considering rainfall as an instrumental variable, without accounting for the implication that these many hypothesized causal pathways would, if taken seriously, represent violations of the assumption of the exclusion restriction. Even if a particular experiment or observational study analyzes only one causal factor, the existence of potential interactions (as indeed are implied if one were to take the social science literature at face value) destroys the implicit assumption that an effect measured under some particular set of conditions can be interpreted as a general or persistent effect.
@@ 2 Piranhas and butterflies
“A butterfly flapping its wings usually has no effect, even in sensitive or chaotic systems. You might even say especially in sensitive or chaotic systems. Sensitive systems are not always and everywhere sensitive to everything. They are sensitive in particular ways under particular circumstances and can otherwise be resistant to influence. …The lesson that many people draw from their first exposure to complex systems is that there are high-leverage points, if only you can find them and manipulate them. They want to insert a butterfly tat just the right time and place to bring about a desired outcome. Instead, we should humbly evaluate to what extent it is possible to steer complex systems at all. We should evaluate what aspects can be steered and how well they can be steered. The most effective intervention may not come from tweaking the inputs but from changing the structure of the system.”
Whether thinking in terms of butterflies or piranhas, we can think of an infinite series of potential effects, which implies that only a few can be large and also creates the possibility of interactions that, after some point, overwhelm any main effects in the population, which in social science corresponds to variation across people and over time (as interventions often have time-limited effects) and in our mathematical models are represented by probability distributions.
@@ 3 Example: hypothesized effect sizes in social priming
We demonstrate the possibility of quantitative analysis of the piranha problem using the example of an influential experiment from 1996 in which participants were given a scrambled-sentence task and then were surreptitiously timed when walking away from the lab.
@@
All these options have major implications for the study of social priming and, more generally, for causal inference in an open-ended setting with large numbers of potential influences. If large interactions are possible, this suggests that stable individual treatment effects might be impossible to find: a 13% effect of a particular intervention in one particular experiment might be $-18\%$ in another context or $+2\%$ in the presence of some other unnoticed factor, and this in turn raises questions about the relevance of any particular study. If effects are much smaller than reported, this suggests that existing studies are extremely underpowered, so that published estimates are drastically overestimated and often in the wrong direction (Gelman and Carlin, 2014), thus essentially noise. At the same time, a restriction of the universe of potential stimuli would require an overhaul of the underlying theoretical framework in which just about any stimulus can cause a noticeable change. For example, if we think there cannot be more than five or ten large effects on walking speed, it would seem a stretch that unnoticed words in a sentence scrambling test would be one of these special factors. Finally, if the distribution of average effects is represented by a long series, most of whose elements are tiny (either because of small individual effects or because any large effects occur infrequently in the general population, as with rare diseases or short-term interventions with rapidly decaying effects), this implies a prior distribution on average effect sizes with a spike near zero, which in turn would result in near-zero estimated population effect sizes in most cases. Our point is not that all effects are zero but rather that in a world of essentially infinite possible causal factors, some external structure must be applied for stable effects to be identified from finite samples.

### v5

@@
The piranha problem: Large effects swimming in a small pond [^1] To appear in the Notices of the American Mathematical Society. We thank Lauren Kennedy, the editor, and four reviewers for helpful discussion. CT and DH acknowledge NSF grant CCF-1740833, a JP Morgan Faculty Award, and a Sloan Research Fellowship. Much of this work was done while CT was at Columbia University. AG acknowledges ONR grant N000142212648. BG acknowledges NSF grants 2051246 and 2153019. AV acknowledges Research Council of Finland project 340721. Affiliations: Memorial Sloan Kettering Cancer Center (CT), Department of Statistics, Columbia University (PG and AG), Department of Political Science, Columbia University (BG and AG), Department of Computer Science, Aalto University (AV), Department of Computer Science, Columbia University (DH).
@@ 1 Background
The implication of the published and well-publicized claims regarding ovulation and voting, shark attacks and voting, college football and voting, etc., is not merely that some voters are superficial and fickle. No, this literature claims that seemingly trivial or irrelevant factors have large and consistent effects, and this runs into the problem of interactions. For example, the effect on your vote of the local college football team losing could depend crucially on whether there’s been a shark attack lately, or on what’s up with your hormones on election day. Or the effect could be positive in an election with a female candidate and negative in an election with a male candidate. Or the effect could interact with your parents’ socioeconomic status, or whether your child is a boy or a girl, or the latest campaign ad, or any of the many other factors that have been studied in the evolutionary psychology and political psychology literatures. If such effects are large, it is necessary to consider their interactions, because the average effect of a factor in any particular study will depend on the levels of all the other factors in that environment. Similarly, Mellon (2020) has argued against naive assumptions of causal identification in economics, where there is a large literature considering rainfall as an instrumental variable, without accounting for the implication that these many hypothesized causal pathways would, if taken seriously, represent violations of the assumptions of the model. Even if a particular experiment or observational study analyzes only one causal factor, the existence of potential interactions (as indeed are implied if one were to take the social science literature at face value) destroys the implicit assumption that an effect measured under some particular set of conditions can be interpreted as a general or persistent effect.
@@ 2 Piranhas and butterflies
“A butterfly flapping its wings usually has no effect, even in sensitive or chaotic systems. You might even say especially in sensitive or chaotic systems. Sensitive systems are not always and everywhere sensitive to everything. They are sensitive in particular ways under particular circumstances and can otherwise be resistant to influence. …The lesson that many people draw from their first exposure to complex systems is that there are high-leverage points, if only you can find them and manipulate them. They want to insert a butterfly at just the right time and place to bring about a desired outcome. Instead, we should humbly evaluate to what extent it is possible to steer complex systems at all. We should evaluate what aspects can be steered and how well they can be steered. The most effective intervention may not come from tweaking the inputs but from changing the structure of the system.”
Effects in social science vary across people and scenarios and over time, and they can be represented by probability distributions. Cook’s advice to think about “the structure of the system” echoes recommendations from the literature on statistical quality control that system-level variation puts a limit on what can be learned about the average effects of particular interventions. In the presence of possible interactions, there is no reason to expect stability of treatment effects.
@@ 3 Example: hypothesized effect sizes in social priming
We demonstrate the possibility of quantitative analysis of the piranha problem using the example of an influential experiment from 1996 in which participants were given a scrambled-sentence task and then were surreptitiously timed when walking away from the lab. Students whose sentences included elderly-related words such as “worried,” “Florida,” “old,” and “lonely” walked an average of 13% more slowly than students in the control condition, and the difference was statistically significant.
@@
All these options have major implications for the study of social priming and, more generally, for causal inference in an open-ended setting with large numbers of potential influences. If large interactions are possible, this suggests that stable individual treatment effects might be impossible to find: a 13% effect of a particular intervention in one particular experiment might be $-18\%$ in another context or $+2\%$ in the presence of some other unnoticed factor, and this in turn raises questions about the relevance of any particular study. If effects are much smaller than reported, this suggests that existing studies are extremely underpowered, so that published estimates are drastically overestimated and often in the wrong direction (Gelman and Carlin, 2014), thus essentially noise. At the same time, a restriction of the universe of potential stimuli would require an overhaul of the underlying theoretical framework in which just about any stimulus can cause a noticeable change. For example, if we think there cannot be more than five or ten large effects on walking speed, it would seem a stretch that unnoticed words in a sentence scrambling test would be one of these special factors. Finally, if the distribution of average effects is represented by a long series, most of whose elements are tiny (either because of small individual effects or because any large effects occur infrequently in the general population, as with rare diseases or short-term interventions with rapidly decaying effects), this implies a prior distribution on average effect sizes with a spike near zero, which in turn would result in near-zero estimated population effect sizes in most cases. Our point is not that all effects are zero but rather that in a world of essentially infinitely many possible causal factors, some external structure must be applied in order to identify stable effects from finite samples.